A mixture of 4% disinfectant solution is to be made from 7% and 2% disinfectant solutions. How much of each solution solution be used of 40 gallons of 4% solution are needed

Answer by addingup(3677)

let the quantity of 7% solution be x
Then, the quantity of 2% will be 40 - x
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
0.07x + 0.02(40-x) = 0.04(40)
0.07x + 0.8 - 0.02x = 1.6
0.05x = 0.8
x = 16 this is the quantity needed of the 7% solution
to be mixed with:
40-16 = 24 gallons of the 2% solution
Happy learning!